OpenStudy (anonymous):

a group of hikers buy a bag of trail mix each bag contains 3 1/2 cups of trail mix and trail mix is shared evenly among 12 hikers how many cups of trail mix with each hiker receive

OpenStudy (anonymous):

(3 1/2) / 12 = (7/2) / 12 = 7/2 * 1/12 = 7/24 cups check.. 7/24 * 12 = 3 1/2 84/24 = 3 1/2 42/12 = 7/2 7/2 = 7/2 (correct) each person gets 7/24 cups